# The hybrid matching of Hurwitz systems

### v1

@@ 1. Introduction
Let $q\in\mathbb{R}^{2}$. If $q\in\Sigma^{\pm}\setminus\Sigma$, then the local trajectory of $\mathcal{X}$ at $q$ is given by the local trajectory of $X^{\pm}$ at $q$. If $q\in\Sigma$, then the local trajectory can be classified as crossing, sliding or escaping. Moreover, $q$ can also be a new type of singularity, known as tangential singularity. In this paper we only deal with crossing points. For the definition of the other types of trajectories and for more details on piecewise smooth vector fields, we refer to [11] and the references therein.
@@
(2) $X^{+}(q),\nabla h(q)\right>\cdot\left<X^{-}(q),\nabla h(q)\right>>0,$
[…]
As far as we know the first approach on the characterization of global asymptotic stability for piecewise smooth vector fields is due Freire to et al [14], where they proved that the continuous matching of two Hurwitz vector fields is GAS. More precisely, if $\mathcal{X}=(X^{+},X^{-};\Sigma)$ is a planar piecewise linear vector field satisfying the following hypotheses:
@@ 3. $(H_{2}^{\prime})$
imply the global asymptotically stability for $\mathcal{X}$.
@@ 2. Hybrid systems
At this instant, the ball bounces back upwards and its velocity undergoes an instantaneous change $v(t_{0})\mapsto-\rho v(t_{0})$, modeled by an inelastic collision with dissipate factor $10<\rho<$ 1. That is, at $h=0$ our system undergoes a jump (or reset) $\varphi$ given by $\varphi(0,v)=(0,-\rho v)$. Roughly speaking, the reset map represents the instantaneous loss of energy that the system suffers when hitting the ground. Observe that any orbit converges to the origin, which represents a ball standing still in the ground.
For more details on the bouncing ball model, we refer to [30, Section $2.2.3$] and [6, Section $1.2$]. For a different example modeling a type of pinball machine with curved surface, we refer to [29, Example $2$]. For a recent survey in the field, we refer to [5].
The notion of a hybrid dynamical system is relatively new. Following Schaft and Schumacher [30, Section $1.1$], we quote:
[$\dots$] the area of hybrid systems is still in its infancy and a general theory of hybrid systems seems premature. More inherently, hybrid systems is such a wide notion that sticking to a single definition shall be too restrictive. [$\dots$] Another difficulty in discussing hybrid systems is that various scientific communities with their own approaches have contributed (and are still contributing) to the area. At least the following three communities can be distinguished. [The] computer science community [$\dots$] [the] modeling and simulation community [$\dots$] [and the] systems and control community.
@@
As anticipated in the Introduction, in this paper we will embedded the the piecewise smooth vector field $\mathcal{X}=(X^{+},X^{-};\Sigma$ in a hybrid structure. This embedding will allow us to draw a precise sufficient and necessary condition for global stability and moreover will point the way for the bifurcation of a limit cycle that would not be possible without the hybrid framework.
@@
(5) $X^{+}(q),\nabla h_{\rho}(q)\right>\cdot\left<X^{-}(q),\nabla h_{\rho}(q)\right>>0,$
for every $q\in\Sigma_{\rho}\setminus\{(0,0)\}$. In particular we recall that the geometric interpretation of (5) is that $X^{\pm}(q)$ are transversal to $\Sigma$ at $q$ and that both vectors $X^{+}(q)$ and $X^{-}(q)$ points to the same direction relatively to $\nabla h_{\rho}(q)$. For hybrid systems without the crossing property and thus allowing other dynamics such as sliding (or sticking), we refer to [30, Section $6.2.3$] and [6, p. $82$].
@@
(6) $\varphi_{\rho}(x,y)=\left\{\begin{array}[]{ll}(-a|x|^{r},0),&\text{if }x\leqslant 0,\vspace{0.2cm}\\ (bx^{s},\rho bx^{s}),&\text{if }x\geqslant 0,\end{array}\right.$
@@
(7) $\varphi_{\rho}^{-1}(x,y)=\left\{\begin{array}[]{ll}\bigl{(}-\bigl{|}a^{-1}x\bigr{|}^{\frac{1}{r}},0\bigr{)},&\text{if }x\leqslant 0,\vspace{0.2cm}\\ \bigl{(}b^{-\frac{1}{s}}x^{\frac{1}{s}},\rho b^{-\frac{1}{s}}x^{\frac{1}{s}}\bigr{)},&\text{if }x\geqslant 0,\end{array}\right.$
@@
The dynamics of $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ works as follows. Given $q_{0}\in\mathbb{R}^{2}\setminus\Sigma$, the local trajectory of $\mathfrak{X}$ at $q_{0}$ is given by the local trajectory of $X^{\pm}$. If this trajectory never intersects $\Sigma_{\rho}$, then we are done. If it does intersect $\Sigma$ at a point $q_{1}$, then we apply the jump obtaining a point $q_{2}=\varphi_{\rho}(q_{1})$. It follows from (5) that the local trajectories of $X^{\pm}$ at $q_{2}$ are transversal to $\Sigma$ and agree on orientation. Hence we are able to follow exactly one of such trajectories leaving $\Sigma$. The process now repeats. See Figure 2.
@@ 4.1. Global normal form
$B^{\pm}=\left(\begin{array}[]{cc}b_{11}^{\pm}&b_{12}^{\pm}\vspace{0.2cm}\\ b_{21}^{\pm}&b_{22}^{\pm}\end{array}\right),$
@@
$C^{\pm}=\left(\begin{array}[]{cc}c_{11}^{\pm}&c_{12}^{\pm}\vspace{0.2cm}\\ c_{21}^{\pm}&c_{22}^{\pm}\end{array}\right),$
@@
$\begin{array}[]{cc}\displaystyle c_{11}^{\pm}=1-\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}}\rho,&\displaystyle c_{12}^{\pm}=\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}},\vspace{0.2cm}\\ \displaystyle c_{21}^{\pm}=\rho+\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}\rho,&\displaystyle c_{22}^{\pm}=-\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}.\end{array}$
@@
$H(x,y)=\left\{\begin{array}[]{ll}C^{+}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{+},\vspace{0.2cm}\\ C^{-}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{-},\end{array}\right.$
@@ 4.2. Hurwitz vector fields of type $N_{1}$
$\begin{array}[]{l}R_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r_{1}^{+}y\},\vspace{0.2cm}\\ E_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{1}^{+}y\},\vspace{0.2cm}\\ R_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon r_{2}^{+}y<x<r_{1}^{+}y\},\vspace{0.2cm}\\ E_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{2}^{+}y\},\vspace{0.2cm}\\ R_{3}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r_{2}^{+}y\},\end{array}$
@@ 4.3. Hurwitz vector fields of type $N_{2}$
$\begin{array}[]{l}S_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r^{+}y\},\vspace{0.2cm}\\ E^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r^{+}y\},\vspace{0.2cm}\\ S_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r^{+}y\},\end{array}$
@@ 4.4. Hurwitz vector fields of type $F$
To end this section, we observe that since $X^{+}$ is linear we can compute explicitly its solution, see [28, Chapter $1$]. Therefore it is not hard to see that if $X^{+}$ is a Hurwitz vector field of type $F$ with the normal form given by Lemma 1, then its solution
@@
(9) $\begin{array}[]{l}\displaystyle x^{+}(t;x,y)=e^{\lambda^{+}t}\left(x\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(\lambda^{+}x-r_{1}^{+}r_{2}^{+}y)\sin(\mu^{+}t)\right),\vspace{0.2cm}\\ \displaystyle y^{+}(t;x,y)=e^{\lambda^{+}t}\left(y\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(x-\lambda^{+}y)\sin(\mu^{+}t)\right).\end{array}$
@@ 5. Proof of Theorem A
From (14) we can deduce statements $(i)$, $(ii)$, $(iii)$ and $(iv)$. More precisely, if $r=1/s$ then we can factor $x$ out of (14) and thus we obtain that the origin is globally asymptotically stable (resp. globally asymptotically unstable) if (14) is satisfied (resp. is the opposite inequality is satisfied). If the equality holds, then the origin is a global center.
@@
Observe that if $a=b=r=s=1$, then the above inequality is satisfied regardless of $\lambda^{\pm}$ and $\mu^{\pm}$, i.e. regardless of $X^{\pm}$. This in addition with Theorem A$(a)$ ensures that if the jump $\varphi_{\rho}$ is reduced to the identity map (i.e. if the hybrid system $\mathfrak{X}$ is reduced to a piecewise linear vector field) and $\Sigma_{\rho}$ is reduced to a straight line, then $\mathfrak{X}$ is GAS. This is precisely the main result of [9].
@@
$\begin{array}[]{ll}\displaystyle\cos(x+\pi)=-\cos x,&\displaystyle\sin(x+\pi)=-\sin(x),\vspace{0.2cm}\\ \displaystyle\cos(\arctan x)=\frac{1}{\sqrt{1+x^{2}}},&\displaystyle\sin(\arctan x)=\frac{x}{\sqrt{1+x^{2}}},\end{array}$
@@
(16) $\lim\limits_{\rho\to+\infty}\frac{|\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})|}{|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}}=\left\{\begin{array}[]{ll}0,&\text{if }r>1,\vspace{0.2cm}\\ +\infty,&\text{if }r<1,\vspace{0.2cm}\\ \displaystyle\sqrt{\frac{(\lambda^{-})^{2}+(\mu^{-})^{2}}{(\lambda^{+})^{2}+(\mu^{+})^{2}}},&\text{if }r=1.\end{array}\right.$
@@
$\begin{array}[]{l}X^{\pm}(x,y)=(-2x-2y,x),\vspace{0.2cm}\\ \Sigma=\{(x,y)\in\mathbb{R}^{2}\colon y=0\},\vspace{0.2cm}\\ \varphi(x,0)=(x^{3},0).\end{array}$
@@ Acknowledgments
The first author is partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais, Brazil, grant APQ-02153-23; by São Paulo Research Foundation (FAPESP), Brazil, grant 2019/07316–0; and by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, grant 311921/2020-5.

### v2

@@ 1. Introduction
Let $q\in\mathbb{R}^{2}$. If $q\in\Sigma^{\pm}\setminus\Sigma$, then the local trajectory of $\mathcal{X}$ at $q$ is given by the local trajectory of $X^{\pm}$ at $q$. If $q\in\Sigma$, then the local trajectory can be classified as crossing, sliding or escaping. Moreover, $q$ can also be a new type of singularity, such as pseudo-singularity, or tangential singularity. See  Guardia et al [18]. In this paper we only deal with crossing points. For the definition of the other types of trajectories and for more details on piecewise smooth vector fields, we refer to [11] and the references therein.
@@
(2) $\left<X^{+}(q),\nabla h(q)\right>\cdot\left<X^{-}(q),\nabla h(q)\right>>0,$
where $\left$$\cdot$$\cdot\right$> denotes the standard inner product of $\mathbb{R}^{2}$. Geometrically (2) means that $X^{+}(q)$ and $X^{-}(q)$ are transversal to $\Sigma$ at $q$ and both point towards either $\Sigma^{+}$ or $\Sigma^{-}$. In particular, the local trajectories of $X^{\pm}$ at $q$ agree on orientation and thus we can define the local trajectory of $\mathcal{X}$ at $q$ as the concatenation to the local trajectories of $X^{\pm}$.
As far as we know the first approach on the characterization of global asymptotic stability for piecewise smooth vector fields is due to Freire et al [14], where they proved that the continuous matching of two Hurwitz vector fields is GAS. More precisely, if $\mathcal{X}=(X^{+},X^{-};\Sigma)$ is a planar piecewise linear vector field satisfying the following hypotheses:
@@ 3. $(H_{2}^{\prime})$
imply the global asymptotically stability for $\mathcal{X}$. We remark that to know whether a piecewise linear system (and in particular piecewise Hurwitz system) is GAS has applicability in control theory. See Iwatani and Hara [21].
@@ 2. Hybrid systems
At this instant, the ball bounces back upwards and its velocity undergoes an instantaneous change $v(t_{0})\mapsto-\rho v(t_{0})$, modeled by an inelastic collision with dissipate factor $0<\rho<1$. That is, at $h=0$ our system undergoes a jump (or reset) $\varphi$ given by $\varphi(0,v)=(0,-\rho v)$. Roughly speaking, the reset map represents the instantaneous loss of energy that the system suffers when hitting the ground. Observe that any orbit converges to the origin, which represents a ball standing still in the ground.
For more details on the bouncing ball model, we refer to [32, Section $2.2.3$] and [6, Section $1.2$]. For a different example modeling a type of pinball machine with curved surface, we refer to [31, Example $2$]. For a recent survey in the field, we refer to [5].
The notion of a hybrid dynamical system is relatively new. Following Schaft and Schumacher [32, Section $1.1$], we quote:
$\dots$] the area of hybrid systems is still in its infancy and a general theory of hybrid systems seems premature. More inherently, hybrid systems is such a wide notion that sticking to a single definition shall be too restrictive. $\dots$] Another difficulty in discussing hybrid systems is that various scientific communities with their own approaches have contributed (and are still contributing) to the area. At least the following three communities can be distinguished. [The] computer science community $\dots$] [the] modeling and simulation community $\dots$] [and the] systems and control community.
@@
As anticipated in the Introduction, in this paper we will embedded the the piecewise smooth vector field $\mathcal{X}=(X^{+},X^{-};\Sigma)$ in a hybrid structure. This embedding will allow us to draw a precise sufficient and necessary condition for global stability and moreover will point the way for the bifurcation of a limit cycle that would not be possible without the hybrid framework.
@@
(5) $\left<X^{+}(q),\nabla h_{\rho}(q)\right>\cdot\left<X^{-}(q),\nabla h_{\rho}(q)\right>>0,$
for every $q\in\Sigma_{\rho}\setminus\{(0,0)\}$. In particular we recall that the geometric interpretation of (5) is that $X^{\pm}(q)$ are transversal to $\Sigma$ at $q$ and that both vectors $X^{+}(q)$ and $X^{-}(q)$ points to the same direction relatively to $\nabla h_{\rho}(q)$. For hybrid systems without the crossing property and thus allowing other dynamics such as sliding (or sticking), we refer to [32, Section $6.2.3$] and [6, p. $82$].
@@
(6) $\varphi_{\rho}(x,y)=\left\{\begin{array}[]{ll}(-a|x|^{r},0),&\text{if }x\leqslant 0,\vskip 5.69046pt\\ (bx^{s},\rho bx^{s}),&\text{if }x\geqslant 0,\end{array}\right$.
@@
(7) $\varphi_{\rho}^{-1}(x,y)=\left\{\begin{array}[]{ll}\bigl(-\bigl|a^{-1}x\bigr|^{\frac{1}{r}},0\bigr),&\text{if }x\leqslant 0,\vskip 5.69046pt\\ \bigl(b^{-\frac{1}{s}}x^{\frac{1}{s}},\rho b^{-\frac{1}{s}}x^{\frac{1}{s}}\bigr),&\text{if }x\geqslant 0,\end{array}\right$.
@@
The dynamics of $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ work as follows. Given $q_{0}\in\mathbb{R}^{2}\setminus\Sigma$, the local trajectory of $\mathfrak{X}$ at $q_{0}$ is given by the local trajectory of $X^{\pm}$. If this trajectory never intersects $\Sigma_{\rho}$, then we are done. If it does intersect $\Sigma$ at a point $q_{1}$, then we apply the jump obtaining a point $q_{2}=\varphi_{\rho}(q_{1})$. It follows from (5) that the local trajectories of $X^{\pm}$ at $q_{2}$ are transversal to $\Sigma$ and agree on orientation. Hence we are able to follow exactly one of such trajectories leaving $\Sigma$. The process now repeats. See Figure 2.
@@ 4.1. Global normal form
$B^{\pm}=\left(\begin{array}[]{cc}b_{11}^{\pm}&b_{12}^{\pm}\vskip 5.69046pt\\ b_{21}^{\pm}&b_{22}^{\pm}\end{array}\right),$
@@
$C^{\pm}=\left(\begin{array}[]{cc}c_{11}^{\pm}&c_{12}^{\pm}\vskip 5.69046pt\\ c_{21}^{\pm}&c_{22}^{\pm}\end{array}\right),$
@@
$\begin{array}[]{cc}\displaystyle c_{11}^{\pm}=1-\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}}\rho,&\displaystyle c_{12}^{\pm}=\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}},\vskip 5.69046pt\\ \displaystyle c_{21}^{\pm}=\rho+\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}\rho,&\displaystyle c_{22}^{\pm}=-\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}.\end{array}$
@@
$H(x,y)=\left\{\begin{array}[]{ll}C^{+}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{+},\vskip 5.69046pt\\ C^{-}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{-},\end{array}\right$.
@@ 4.2. Hurwitz vector fields of type $N_{1}$
$\begin{array}[]{l}R_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r_{1}^{+}y\},\vskip 5.69046pt\\ E_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{1}^{+}y\},\vskip 5.69046pt\\ R_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon r_{2}^{+}y<x<r_{1}^{+}y\},\vskip 5.69046pt\\ E_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{2}^{+}y\},\vskip 5.69046pt\\ R_{3}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r_{2}^{+}y\},\end{array}$
@@ 4.3. Hurwitz vector fields of type $N_{2}$
$\begin{array}[]{l}S_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r^{+}y\},\vskip 5.69046pt\\ E^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r^{+}y\},\vskip 5.69046pt\\ S_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r^{+}y\},\end{array}$
@@
To end this section, we observe that since $X^{+}$ is linear we can compute explicitly its solution, see [30, Chapter 11]. Therefore it is not hard to see that if $X^{+}$ is a Hurwitz vector field of type $F$ with the normal form given by Lemma 1, then its solution
@@
(9) $\begin{array}[]{l}\displaystyle x^{+}(t;x,y)=e^{\lambda^{+}t}\left(x\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(\lambda^{+}x-r_{1}^{+}r_{2}^{+}y)\sin(\mu^{+}t)\right),\vskip 5.69046pt\\ \displaystyle y^{+}(t;x,y)=e^{\lambda^{+}t}\left(y\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(x-\lambda^{+}y)\sin(\mu^{+}t)\right).\end{array}$
@@ 5. Proof of Theorem A
From (14) we can deduce statements $(i)$, $(ii)$, $(iii)$ and $(iv)$. More precisely, if $r=1/s$ then we can factor $x$ out of (14) and thus obtain that the origin is globally asymptotically stable (resp. globally asymptotically unstable) if (14) is satisfied (resp. is the opposite inequality is satisfied). If the equality holds, then the origin is a global center.
@@
Observe that if $a=b=r=s=1$, then the above inequality is satisfied regardless of $\lambda^{\pm}$ and $\mu^{\pm}$, i.e. regardless of $X^{\pm}$. This in addition with Theorem A(a)(a) ensures that if the jump $\varphi_{\rho}$ is reduced to the identity map (i.e. if the hybrid system $\mathfrak{X}$ is reduced to a piecewise linear vector field) and $\Sigma_{\rho}$ is reduced to a straight line, then $\mathfrak{X}$ is GAS. This is precisely the main result of [9].
@@
$\begin{array}[]{ll}\displaystyle\cos(x+\pi)=-\cos x,&\displaystyle\sin(x+\pi)=-\sin(x),\vskip 5.69046pt\\ \displaystyle\cos(\arctan x)=\frac{1}{\sqrt{1+x^{2}}},&\displaystyle\sin(\arctan x)=\frac{x}{\sqrt{1+x^{2}}},\end{array}$
@@
(16) $\lim\limits_{\rho\to+\infty}\frac{|\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})|}{|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}}=\left\{\begin{array}[]{ll}0,&\text{if }r>1,\vskip 5.69046pt\\ +\infty,&\text{if }r<1,\vskip 5.69046pt\\ \displaystyle\sqrt{\frac{(\lambda^{-})^{2}+(\mu^{-})^{2}}{(\lambda^{+})^{2}+(\mu^{+})^{2}}},&\text{if }r=1.\end{array}\right$.
@@
$\begin{array}[]{l}X^{\pm}(x,y)=(-2x-2y,x),\vskip 5.69046pt\\ \Sigma=\{(x,y)\in\mathbb{R}^{2}\colon y=0\},\vskip 5.69046pt\\ \varphi(x,0)=(x^{3},0).\end{array}$
@@ Acknowledgments
The first author is partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais, Brazil, grant numbers APQ-02153-23, APQ-05207-23 and RED-00133-21.